OpenStudy (halfdeafdarling):

Need quick help on a functions problem, anyone able to assist? I am a bit confused because I completed the rest of the problems like this but no matter what I have done it baffles me! ~The number of hours (H) that a candle will burn increases when the length of the candle (L) increases. Write the correct equation for this scenario, and solve for the number of hours when the length is 2. This is the question the the table it length hours 15 3 20 4 I have already done the set up H=?/L its just every time I plug in numbers I get a crazy answer that doesn't fit

random231 (random231):

if H increses, so does L which means H is directly proportional to L

OpenStudy (halfdeafdarling):

would my answer need to be h=45/l= 22.5 ? 45/2 = 22.5 and doing the other equations get 45 as the answer when doing so, would that make sense for it to be 22.5? Feel like I am going insane.

random231 (random231):

nuh, first you need to know what directly proportional means.

random231 (random231):

 where k is a constant